A 0.582-mil sample of an ideal gas is at 15degrees Celsius and 81.8kPa. Calculate the volume for the ideal gas as described by the ideal gas law
1
Expert's answer
2020-05-27T13:34:31-0400

Solution:

The first step is to convert temperature to Kelvins and pressure to atm:

15°C + 273 = 288 K

81.8 kPa × (1 atm / 101.325 kPa) = 0.8073 atm


The Ideal Gas Law:

PV = nRT

Now we can substitute the conditions into the ideal gas law:

(0.8073 atm) × (V L) = (0.582 mol) × (0.08206 L atm mol-1 K-1) × (288 K)

(0.8073 atm) × (V L) = (13.7546 L atm)

V = 17.04 L


Answer: V = 17.04 L